OpenStudy (anonymous):

The formula for degrees Celsius is: C 5/9 = (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is. The formula for degrees Celsius is: C 5/9 = (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is. @Mathematics

OpenStudy (anonymous):

(1) F=C(5/9)-32 (2) F=10(5/9)-32=338/9(=37.5..)